0.9m Telescope "Issues"


   DATE         EVENT                                                      STATUS

15. MAR 2017   images don't write during sequence                          .....
               might happen when increasing the # exposures	         
14. MAR 2017   test frames saved as normal frames                          .....
               when the test button is clicked, the frame is named
               with the input character string, seems to happen when
               editing file title

13. FEB 2017   test image numbers do not increment properly                .....

               If you reset Image Number = 1, then take 1 test zero
               and 1 test bias for calibrations, the Image Number
               moves to 5, when it should still read 1 (because they
               are test frames).
12. FEB 2017   automatically name output of focus exposure to "focus"      .....

               when focus sub-exposures are taken, the Image Number
               goes up by one for every sub-exposure and the final
               focus exposure is saved as the file with name given in
               "Basename" field and with the number of the last Image
               Number. It would be nice to change the focus image name
               to e.g. "focus" in the same manner as it is done with
               TEST exposure, and leave Image Number unchanged.

11. JAN 2017   windows machine cannot connect to BIW computer              .....

               known problem, need smooth way out (Marco delivered?)

               error message is "an operation on a socked could not be
               performed because the system lacked sufficient buffer
               space. Do you wish to attempt to reconnect to
10. DEC 2016   remaining trails to west on bright stars                    .....

               from Marco 2016.0718: "Peter can give you more details
               on this, but basically they are building a new Torrent
               system (got all the parts in house now) that will be
               coupled to the spare 0.9m in the lab. With that setup
               it should be much simpler/faster/efficient to follow up
               the trails. Peter and the new EE (Braulio) are also
               designing a new interface board that may be much better
               in terms of trails and noise. I would say that if the
               science is not being affected, wait for lab results
               before doing anything else. Peter should be able to
               give you more details and timing on this."

09. DEC 2016   read noise 2X higher than original 8-10e-                   .....

               2017.0227 Peter and Braulio reduce noise to 15 e-

               2017.1201 Jonathan Irwin found read noise = 19.158 e- and
               gain = 3.046

08. DEC 2016   have white "dofocus" window stay open and have message       OK?
               "The double shift indicates the end of the sequence."

07. DEC 2016   TCS-TIME in headers incorrect date and wrong format         .....

               2017.0315 format ok, but date off by 2 days

               The TCS displays the correct UT and date, but BIW
               displays the correct UT and a date that is 1 day
               behind.  This incorrect date is written into the
               headers for TCS-TIME and in the reversed format:
               TCS-TIME   2016-06-12  for UT 07 December (incorrect)
               DATE-OBS   2016-12-07  for UT 07 December (correct)

06. DEC 2016   change XPIXSIZE and YPIXSIZE to 0.401" in headers           FIXED

05. DEC 2016   tweak BIW GUI buttons                                       FIXED

04. SEP 2016   incorrect #exposures sometimes taken                        .....
03. SEP 2016   hour angles in headers for east frames positive             FIXED
02. SEP 2016   test frame not written to current directory                 FIXED
01. MAY 2016   bad trails to west on bright stars                          BETTER